Given a standard normal distribution with mean =60 and standard deviation = 35, find P(X<75)


1
Expert's answer
2022-01-27T14:43:57-0500
"P(X<75)=P(Z<\dfrac{75-60}{35})"

"=P(Z<\dfrac{3}{7})\approx P(Z<0.42857)"

"\approx0.665882"